40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 6 21:58:39 2013, Jenne, Update, Electronics, POX whitening was fine all along 
    Reply  Wed Mar 6 22:11:19 2013, Koji, Update, Electronics, POX whitening was fine all along 
       Reply  Thu Mar 7 18:41:03 2013, Jenne, Update, Electronics, POX whitening was fine all along POX11I.pngPOX11Q.png
    Reply  Thu Mar 7 11:43:15 2013, Jenne, Update, Electronics, POX and POY whitening DC gain left low 
Message ID: 8246     Entry time: Wed Mar 6 21:58:39 2013     Reply to this: 8247   8249
Author: Jenne 
Type: Update 
Category: Electronics 
Subject: POX whitening was fine all along 

After my investigations this afternoon (with help from Sendhil and Shivaraj), I do not find any problems with the POX whitening switching.

Earlier this afternoon / evening I was misleading myself into thinking that either the switching component (ADG333ABR) was broken, or that the whitening op amps (LT1124CS8) were broken on the POX I&Q and POY I&Q channels.  I had not realized until Jamie mentioned the possibility, that some of the DC gain stages were on for POX and POY.  POX and POY (I&Q for both) all had +36dB of gain, so when I was injecting my 60Hz sine wave into those channels, the whitening opamps were already saturated, which is why it didn't look like I was getting any gain.  When I set them all to 0dB (which is what AS11 and REFL11, the other 2 PDs using that whitening board, were set to), all 8 channels behaved the same.

The shaped whitening (which is either bypassed or not, depending on the condition of the software "unwhite" switch) is 2 filters in series, each with a zero at 15Hz, and a pole at 150Hz, with DC gain of 0dB.  For a 60Hz sine wave, this gives a factor of ~4 from each stage.  After setting all of the whitening gains to 0dB, I was able to see on all 8 channels of the board an input sine wave, a larger (by 4-ish) sine wave, and then a larger (by 4ish again) sine.  When I looked at the output of the switch of all 8 channels, the signal was either the same as the input amplitude, or the same as after the 2nd whitening stage, depending on the "unwhite" filters.

Before looking at actual signals, Sendhil and I also had checked to see that indeed, the board was receiving the digital signal input to the switch chip, requesting switching based on the state of the "unwhite" filters.

I looked through the elogs, and the only "symptoms" I find are from an IFO check-up session that Koji, Den and I had back in May, where we declared in the elog that POX whitening may or may not be switching.  See elog 6595.  We didn't mention what the actual symptoms we saw were, so unless Koji or Den remember something that I don't, I cannot confirm that we are no longer seeing those symptoms.  However, based on the number of "?" after "POX whitening not toggling the analog whitening", I don't think that we were totally sure that something was wrong in the first place.

Anyhow, the whitening board in the LSC rack labeled "WF1", serving AS11, REFL11, POX11 and POY11 has had a thorough checkup, and I give it a clean bill of health.

ELOG V3.1.3-